The following is total monthly budgeted cost and activity information for the four activity centers in the billing department of Oregon Power Company:

Activity Center

Variable

Fixed

Cost Driver

Account inquiry

$80,916

$151,000

3,300 labor hours

Correspondence

$10,034

$24,000

2,900 letters

Account billing

$122,500

$70,000

2,450,000 lines

Bill verification

$12,760

$76,000

22,000 accounts


In September, actual costs and activity were as follows:

Activity Center

Total Costs

Driver Amount

Account inquiry

$231,540

3,270 labor hours

Correspondence

$33,946

2,830 letters

Account billing

$200,676

2,600,000 lines

Bill verification

$89,467

22,180 accounts


Required
Compute the flexible-budget variances for the following two activity cost items (round all answers to the nearest dollar and enter favorable variances as positive numbers and unfavorable variances as negative numbers):

  Account billing    

  Correspondence   
Answer #1
Flexible-budget
Account billing   200000 =70000+(122500/2450000*2600000)
Correspondence    33792 =24000+(10034/2900*2830)
Flexible-budget variances
Account billing   -676 =200000-200676
Correspondence    -154 =33792-33946